LCD Industries purchased a supply of electronic components from Entel Corporation on November 1, 2021. In payment for the $25.0 million purchase, LCD issued a 1-year installment note to be paid in equal monthly payments at the end of each month. The payments include interest at the rate of 24%.(FV of $1. PV of $1. FVA of $1. PVA of $1. FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.) 


Required:

 1. & 2. Prepare the journal entries for LCD's purchase of the components on November 1, 2021 and the first installment payment on November 30, 2021.

 3. What is the amount of Interest expense that LCD will report in its income statement for the year ended December 31, 2021?
Answer #1

Purchase Price : $25 Million

Period : 12 Months

Monthly Interest Rate : 24% / 12 = 2%

Annuity rate : 1/(1.02)^1 + 1/(1.02)^2 +1/(1.02)^3 +................1/(1.02)^12

= 10.57534

Monthly Installment : $25,000,000 / 10.57534

= $2,363,990

1. Journal Entries :

Date Particulars Debit Credit
November 1, 2021 Purchase $25,000,000
Note Payable $25,000,000
(To record the purchase of electronic components)
November 31, 2021 Note payable $1,863,990
Interest Expense $500,000
Cash $2,363,990
(To record the first installment)

2.

Interest Expense for the year ending december 31, 2021

Purchase price : $25,000,000

Interest for first month : $25,000,000 * 2% = $500,000

Installment : $2,363,990

Carryover value : $25,000,000 + $500,000 - $2,363,990

= $23,136,010

Interest for second month : $23,136,010 * 2% = $462,720

Interest for year ending December 2021 :

: $500,000 + $462,720 = $962,720

You can ask your queries in comment section. Please rate if you found this helpful.
Answer #2

Solution 1&2:

Monthly interest rate = 24%/12 = 2%

Monthly installment amount = $25,000,000 / Cumulative PV factor at 2% for 12 periods

= $25,000,000 / 10.569 (note)

= $2,365,408

Note:-

year

Pv factor (2%)

1

0.98

2

0.961

3

0.942

4

0.923

5

0.905

6

0.887

7

0.87

8

0.853

9

0.836

10

0.82

11

0.804

12

0.788

total

10.569

For 2%=1 * 2%=0.02 then add 1

    =1+.02

=1 / 1.02=0.980(1st period)

Then 2nd =0.980 / 1.02=0.961 calculate like this will get all above values.

Journal Entries - LCD

Date

Particulars

Debit

Credit

1-Nov-18

Equipment Dr

$25,000,000.00

            To Notes Payable

$25,000,000.00

(To record purchase of component by issue of note)

30-Nov-18

Interest expense Dr ($25,000,000*2%)

$500,000.00

Notes payable Dr (2,365,408-500,000)

$1,865,408.00

            To Cash

$2,365,408.00

(To record installment payment)

Solution:-3 Calculation of interest expense for December 31, 2021 is shown below

Interest expense Nov

25000000*2%

$500,000

Interest expense Dec

(25000000-1865408)*2%

$462,692

Interest expense

$962,692

Thus, interest expense of $962,692 would be reported in income statement for year ended 31st December 2021

If you are satisfied with my workings please give rating Thank you....